# In-depth characterization of the Kepler-10 three-planet system with HARPS-N RVs and Kepler TTVs

### v1

@@
In-depth characterization of the Kepler-10 three-planet system
with HARPS-N RVs and Kepler TTVs
@@ 1 Introduction
The determination of accurate and precise masses and bulk densities of transiting small planets with radial-velocity (RV) follow-up allows us to infer their composition, under reasonable assumptions on planet interior differentiation (e.g., Zeng & Seager 2008) or mixing (e.g., Dorn et al. 2017). The ever-increasing number of well-characterized small planets has revealed a surprising variety in their densities, and hence compositions. Generally, high bulk densities correspond to rocky planets with either terrestrial (e.g., Dressing et al. 2015) or iron rich (e.g., Bonomo et al. 2019) interiors. Intermediate densities can be reproduced with massive icy shells (water worlds) and/or thin H/He (hydrogen/helium) atmospheres shrouding the planet rocky interiors (e.g., Mortier et al. 2018), with a strong degeneracy between the icy and gaseous mass fractions (e.g., Rogers & Seager 2010; Spiegel et al. 2014); low densities definitively require thicker H/He atmospheres (e.g., Lissauer et al. 2011).
@@
The fundamental goal in estimating planet compositions lies in understanding the main mechanisms that give rise to the observed compositional properties and the related well-known valley in the distribution of planetary radii around solar-type stars at $R_{\rm trans}\sim 1.7-2.0$ $\mathrm{R}_{\oplus}$ , which likely separates prevalently rocky planets ($R_{\rm p}\lesssim R_{\rm trans}$) from volatile-rich ones ($R_{\rm p}\gtrsim R_{\rm trans}$) (Fulton et al. 2017; Van Eylen et al. 2018). Such mechanisms include planet formation at different regions of the protoplanetary disks – dry or ice-rich regions inside or beyond the water iceline, respectively (Lopez 2017; Zeng et al. 2019) – and/or at different timescales, i.e. in gas-rich or gas-poor environments. Several post-formation processes that may afterwards modify the planet composition have been identified, among which i) atmospheric photoevaporation driven by intense stellar XUV radiation, mainly in the first $\sim 100$ Myr after formation (e.g., Owen & Wu 2013, 2017; Lopez & Fortney 2014); ii) core-powered mass loss, i.e. atmospheric escape due to the luminosity of the cooling rocky core, acting on longer timescales ($\sim 1$Gyr) (Ginzburg et al. 2018; Gupta & Schlichting 2019); and iii) erosion due to giant impacts (e.g., Liu et al. 2015; Reinhardt et al. 2022). The observed diversity in planet compositions and the radius valley are likely caused by more than one mechanism. Nonetheless, if one of the above-mentioned (formation or post-formation) mechanisms dominates over the others, it may produce a distinct slope in the position of the radius valley as a function of orbital period, indicating a possible variation of the transition between rocky and non-rocky planets with the level of incident flux received by the planet (e.g., Lopez & Rice 2018; Martinez et al. 2019). This possible dependence of the rocky/non-rocky transition with orbital period must then be compared with the inferred compositions from bulk densities, which are expected to be rocky or volatile-rich below and above the radius valley, respectively, with the changing level of incident flux. For that purpose, more and more precise measurements of bulk densities of small planets with relatively long orbital periods (hence lower insolation), i.e. $P\gtrsim 30$ d, are needed, as only a few of them have been well characterized to date.
@@
Determining the mass and bulk density of small planets with long orbital periods can be much more difficult than for shorter-period, inner planets, mainly for three reasons: firstly, the semi-amplitude of the (Doppler) RV signal induced by the planet decreases as $P^{-1/3}$; secondly, wider temporal baselines of RV measurements are required to properly sample long-period signals; thirdly, these signals may be more affected by correlated noise due to stellar magnetic activity variations and/or the presence of additional planets in multi-planet systems. A clear example of this difficulty is the planetary system K2-3, which consists of three small planets – K2-3 b, c and d with radii $R_{\rm p}=2.3$, 1.8 and 1.6 $\mathrm{R}_{\oplus}$  – orbiting an early M dwarf with periods $P=10.05$, 24.65, and 44.56 d, the latter in the habitable zone (Crossfield et al. 2015). Even with more than 300 HARPS-N and HARPS high-precision radial velocities, it was not possible to determine the masses/densities of the two outer planets with a precision better than $3\sigma$ (Damasso et al. 2018; see also Kosiarek et al. 2019).
@@
Kepler-10 is a “historical” planetary system containing two transiting planets: Kepler-10 b, the first rocky planet discovered by the Kepler space mission (Batalha et al. 2011), with radius $R_{\rm b}=1.47\pm 0.03$ $\mathrm{R}_{\oplus}$ and ultra-short orbital period (USP) $P_{\rm b}=0.837$ d; and Kepler-10 c, which has both a larger radius $R_{\rm b}=2.35_{-0.04}^{+0.09}$ $\mathrm{R}_{\oplus}$ and a much longer orbital period $P_{\rm c}=45.294$ d (Fressin et al. 2011; Dumusque et al. 2014, hereafter D14). The planets orbit a 10.5-Gyr old GV star, which is thus expected to be magnetically quiet (average CaII activity index $\log{R^{{}^{\prime}}_{\rm HK}}\simeq-4.97$), and is located at a distance of $186$ pc.
@@
While there has been consensus on the mass and density of the inner rocky planet Kepler-10 b, determinations of the mass and density of Kepler-10 c in the literature have been largely discrepant, illustrating that the characterization of small planets with long orbital periods may be considerably problematic, even at low levels of stellar magnetic activity. For instance, with 148 high-precision RVs collected with the HARPS-N spectrograph at the Telescopio Nazionale Galileo (Cosentino et al. 2012, 2014), D14 found a RV semi-amplitude of $K_{\rm c}=3.26\pm 0.36$  m s-1, which implies a planetary mass and density of $M_{\rm c}=17.2\pm 1.9$ $\mathrm{M}_{\oplus}$ and $\rho_{\rm c}=7.1\pm 1.0\leavevmode\nobreak\ \rm g\,cm^{-3}$. By using more sparse 72 RV measurements acquired with the HIRES spectrograph at the Keck telescope (Howard et al. 2010), Weiss et al. (2016) (hereafter W16) found $K_{\rm c}=1.09\pm 0.58$  m s-1, corresponding to $M_{\rm c}=5.7_{-2.9}^{+3.2}$ $\mathrm{M}_{\oplus}$ and $\rho_{\rm c}=2.4_{-1.2}^{+1.4}\leavevmode\nobreak\ \rm g\,cm^{-3}$ (one may actually even wonder whether, with a statistical significance of 1.9 $\sigma$, the signal of Kepler-10 c is truly detected in the HIRES RV data). By combining the 148 HARPS-N and the 72 HIRES RVs, W16 obtained $K_{\rm c}=2.67\pm 0.34$  m s-1, and hence $M_{\rm c}=13.98\pm 1.79$ $\mathrm{M}_{\oplus}$ and $\rho_{\rm c}=5.94\pm 0.75\leavevmode\nobreak\ \rm g\,cm^{-3}$. However, this combination of the HARPS-N and HIRES data did not solve the issue of the evident mass discrepancy (see Fig. 5, bottom right panel in W16), but led to an average mass, which was found to be closer to the value found by D14 because the number of HARPS-N RVs was approximately twice that gathered with HIRES.
@@
Table 1: Radial-velocity semi-amplitudes ($K$) and corresponding masses ($M_{\rm p}$) of the Kepler-10 planets from the literature.
@@
In addition to comparing the mass determinations obtained with HARPS-N and HIRES, W16 confirmed the variations in the Kepler-10 c transit times with an amplitude of $\sim 5$ min, which were previously discovered by Kipping et al. (2015). Such transit timing variations (TTVs) indicate the presence of an additional planet in the system, Kepler-10 d, which dynamically perturbs Kepler-10 c. In an attempt to take the constraints of TTVs into account, W16 performed a preliminary combined fit of TTVs and RVs, though i) considering for simplicity a limited number of possible orbital configurations, near 2:1 or 3:2 resonance, for Kepler-10 d, and ii) fixing the so-called TTV “super-period” at the peak of the periodogram of the TTVs at 475.22 d (see Sect. 6 in W16). By doing so, W16 found as their best solution that Kepler-10 d is in apparent 2:1 resonance with Kepler-10 c, and has a period of $101.36$ d and a mass of 6.84 $\mathrm{M}_{\oplus}$
@@
To try and reconcile the discrepancy in the RV semi-amplitude, and hence planetary mass/density of Kepler-10 c, with the HIRES and HARPS-N data, Rajpaul et al. (2017) employed a framework based on Gaussian Process (GP) regression with a quasi-periodic kernel to account for possible correlated noise (e.g., Haywood et al. 2014; Grunblatt et al. 2015), by simultaneously modeling the RVs and some activity indicators, such as the $\log{R^{{}^{\prime}}_{\rm HK}}$ and bisector activity indices (Rajpaul et al. 2015). As a result they found $K_{\rm c}=1.27^{+0.42}_{-0.35}$  m s-1and $K_{\rm c}=1.64^{+0.42}_{-0.34}$  m s-1from the HIRES and HARPS-N data, respectively, and $K_{\rm c}=1.41^{+0.25}_{-0.23}$  m s-1when combining both datasets. The latter $K_{\rm c}$ would imply $M_{\rm c}=7.37^{+1.32}_{-1.19}$ $\mathrm{M}_{\oplus}$ and $\rho_{\rm c}=3.14^{+0.63}_{-0.55}\leavevmode\nobreak\ \rm g\,cm^{-3}$, which are considerably lower than those measured by both D14 and W16. From model comparison using Bayesian evidences, Rajpaul et al. (2017) also found that the 3-planet model was favored over the 2-planet one, and reported both orbital period and mass of Kepler-10 d, i.e. $P_{\rm d}=102\pm 1$ d and $M_{\rm d}=5.90^{+1.70}_{-1.01}$ $\mathrm{M}_{\oplus}$ , in agreement with the findings of W16. Moreover, they found a higher evidence for the GP model, and derived a period of $55.5\pm 0.8$ d for the quasi-periodic variations, though it is not clear whether that period is related to the stellar rotation (Rajpaul et al. 2017).
@@
Table 1 summarizes the RV semi-amplitudes and corresponding masses of the Kepler-10 planets as retrieved in the different above-mentioned works from the literature. Being clear from all these works that Kepler-10 is a complex planetary system, we continued to monitor it in 2017, 2019 and 2020, by practically doubling the number of HARPS-N RVs within the HARPS-N/GTO program. After extracting all the HARPS-N RVs with upgraded data reduction softwares (Sect. 2.2), we analyzed them with several tools (Sect. 3), also in combination with the Kepler-10 c TTVs (Sect. 3.5). In this way we improved the characterization of the Kepler-10 planetary system (Sect. 4) with regard to both its orbital architecture and the determination of precise masses, and thus possible compositions, of the Kepler-10 b and Kepler-10 c transiting planets. This in turn allowed us to put additional constraints on the formation and evolution of the Kepler-10 planetary system (Sect. 5).
@@ 2.1 Kepler photometry
Even though the Kepler light curve was analyzed in several previous works (Batalha et al. 2011; Fogtmann-Schulz et al. 2014; D14; W16; Dai et al. 2019; Singh et al. 2022), we performed a new analysis of the Kepler data to i) re-compute the twenty-four transit times of Kepler-10 c in short (58 s) cadence, and the first two transit epochs observed in long-cadence mode (29.4 min), which are not provided by W16; and ii) re-derive the Kepler-10 c transit parameters.
@@ 2.2.1 HARPS-N Radial velocities
In total, we gathered 308 HARPS-N spectra, 55 of which before the failure of the red side of the charge-coupled device (CCD) in late September 2012 (Bonomo et al. 2014), and 253 after the replacement of the CCD. The former 55 spectra were reduced with the original HARPS-N Data Reduction Software (DRS) version 3.7, and the latter 253 with the ESPRESSO DRS version 2.3.5 adapted to HARPS-N (Dumusque 2021). The radial velocities were extracted from all the spectra by cross-correlating them with a G2V stellar template (e.g., Pepe et al. 2002). The new DRS pipeline allows for a better long-term RV accuracy, due to a careful selection of non-saturated Thorium and Argon lines that are used to calibrate the spectrograph wavelength solution. In the DRS-3.7 Kepler-10 RV data, significant variations in the flux of the original Thorium-Argon calibration lamp induce an artificial quadratic long-term trend of $\sim 5$  m s-1 till the replacement of the Thorium-Argon lamp at the beginning of June 2020, which further produces an RV offset at the level of a  m s-1. Those systematics are gone or at least strongly mitigated to a level below the  m s-1  when using the new pipeline. In addition, the new DRS uses a novel algorithm to determine the wavelength solution that is more stable from calibration to calibration. The night-to-night RV rms offset due to wavelength solution calibration is reduced from $\sim$80 to $\sim$50 cm s-1(Dumusque 2021). Yet, no significant improvements from the new DRS are expected for the first 55 HARPS-N spectra taken with the old CCD, given both the relatively small amount of data and the short timespan coverage. For this reason, we used the original HARPS-N DRS (version 3.7) for them.
@@
Given the high complexity of the Kepler-10 system, as shown by the discrepancy in the Kepler-10 c mass determination from previous works, we also made use of the YARARA tool (Cretignier et al. 2021, 2022) to correct for telluric lines and subtle instrumental effects which cannot be accounted for by the new DRS pipeline, as well as for possible stellar activity variations, despite the low stellar activity level. In doing so, we had to discard 17 out of 253 HARPS-N spectra ($\sim 7$%) because they did not pass the S/N and vetting criteria of YARARA.
@@
Both the 55 DRS-3.7 and 236 YARARA-v2 HARPS-N RVs used in this work are shown in Fig. 1. They were released by Bonomo et al. (2023) along with the stellar activity indicators Full Width at Half Maximum (FWHM), Contrast and Bisector Span of the CCF, and the CaII H&K S-index and $\log{R^{{}^{\prime}}_{\rm HK}}$ [^1] We point out that the CCF Contrast and FWHM activity indicators are affected by strong variations, which are highly correlated, due to changes of the HARPS-N focus. However, this did not result in any systematic RV variation, as the product $\rm Contrast\cdot FWHM$ is conserved.. The improvement achieved with YARARA can be seen in Figure 2 showing the increase in power of the peaks at the expected orbital periods of both Kepler-10 b and Kepler-10 c (from Kepler photometry) in the Generalized Lomb-Scargle (GLS, Zechmeister & Kürster 2009) periodograms of the 236 HARPS-N RVs reduced with the new DRS (top panel) and the new DRS+YARARA-v2 (bottom panel). For this reason, we used the latter for the 236 RVs collected from 2013 to 2021 (no remarkable improvement is instead expected for the 55 RVs gathered earlier for the same aforementioned reasons that we did not apply the new DRS for their extraction). We note that we checked for possible outliers in the HARPS-N RVs by using the Chauvenet’s criterion (e.g., Bonomo et al. 2023), and found none.
@@
Figure 1: Kepler-10 radial velocities. Filled blue and magenta circles show the HARPS-N data collected with the new and old CCD, respectively, and empty green triangles display the HIRES measurements. Radial-velocity zero points as determined with the DE-MCMC analysis (Sect. 3.3.1 and Table LABEL:table_harpsn-hires_param) were subtracted from each dataset.
@@
Figure 2: Generalized Lomb-Scargle periodograms of the 236 HARPS-N radial velocities as reduced with the new DRS (top panel) and the new DRS+YARARA-v2 (bottom panel). The power of the periodograms was normalised by the 1% false alarm probability (FAP) level. Note the increase in power of the peaks at the periods of Kepler-10 b and Kepler-10 c (vertical green lines) with the YARARA reductions.

### v2

# In-depth characterization of the Kepler-10 three-planet system with HARPS-N radial velocities and Kepler transit timing variations

@@
In-depth characterization of the Kepler-10 three-planet system with
HARPS-N radial velocities and Kepler transit timing variations
@@ 1 Introduction
The determination of accurate and precise masses and bulk densities of transiting small planets with radial-velocity (RV) follow-up allows us to infer their composition, under reasonable assumptions of planet interior differentiation (e.g., Zeng & Seager 2008) or mixing (e.g., Dorn et al. 2017). The ever-increasing number of well-characterized small planets has revealed a surprising variety in their densities, and hence their compositions. Generally, high bulk densities correspond to rocky planets with either terrestrial (e.g., Dressing et al. 2015) or iron-rich (e.g., Bonomo et al. 2019) interiors. Intermediate densities can be reproduced by massive icy shells (water worlds) and/or thin H/He (hydrogen/helium) atmospheres shrouding the planet rocky interiors (e.g., Mortier et al. 2018), with a strong degeneracy between the icy and gaseous mass fractions (e.g., Rogers & Seager 2010; Spiegel et al. 2014); low densities definitively require thicker H/He atmospheres (e.g., Lissauer et al. 2011).
@@
The fundamental goal in estimating planet compositions lies in understanding the main mechanisms that give rise to the observed compositional properties and the related well-known valley in the distribution of planetary radii around solar-type stars at $R_{\rm trans}\sim 1.7-2.0$ $\mathrm{R}_{\oplus}$ , which likely separates prevalently rocky planets ($R_{\rm p}\lesssim R_{\rm trans}$) from volatile-rich ones ($R_{\rm p}\gtrsim R_{\rm trans}$; Fulton et al. 2017; Van Eylen et al. 2018). Such mechanisms include planet formation at different regions of the protoplanetary disks – dry or ice-rich regions inside or beyond the water iceline, respectively (Lopez 2017; Zeng et al. 2019) – and/or at different timescales, that is, in gas-rich or gas-poor environments. Several post-formation processes that may afterwards modify the planet composition have been identified, among which (i) atmospheric photoevaporation driven by intense stellar XUV radiation, mainly in the first $\sim 100$ Myr after formation (e.g., Owen & Wu 2013, 2017; Lopez & Fortney 2014); (ii) core-powered mass loss, that is, atmospheric escape due to the luminosity of the cooling rocky core, acting on longer timescales ($\sim 1$ Gyr; Ginzburg et al. 2018; Gupta & Schlichting 2019); and (iii) erosion due to giant impacts (e.g., Liu et al. 2015; Reinhardt et al. 2022). The observed diversity in planet compositions and the radius valley are probably caused by more than one mechanism. Nonetheless, if one of the abovementioned (formation or post-formation) mechanisms dominates over the others, it may produce a distinct slope in the position of the radius valley as a function of orbital period, which indicates a possible variation of the transition between rocky and non-rocky planets with the level of incident flux received by the planet (e.g., Lopez & Rice 2018; Martinez et al. 2019). This possible dependence of the rocky/non-rocky transition with orbital period must then be compared with the inferred compositions from bulk densities, which are expected to be rocky or volatile-rich below and above the radius valley, respectively, with the changing level of incident flux. For that purpose, more precise measurements of the bulk densities of small planets with relatively long orbital periods (hence lower insolation), $P\gtrsim 30$ d, are needed, as only a few of them have been characterized well to date.
@@
Determining the mass and bulk density of small planets with long orbital periods can be much more difficult than for shorter-period, inner planets, for three main reasons. Firstly, the semiamplitude of the (Doppler) RV signal induced by the planet decreases as $P^{-1/3}$; secondly, wider temporal baselines of RV measurements are required to properly sample long-period signals; and thirdly, these signals may be more affected by correlated noise due to stellar magnetic activity variations and/or the presence of additional planets in multiplanet systems. A clear example of this difficulty is the planetary system K2-3, which consists of three small planets – K2-3 b, c, and d with radii $R_{\rm p}=2.3$, 1.8, and 1.6 $\mathrm{R}_{\oplus}$  – which orbit an early M dwarf with periods $P=10.05$, 24.65, and 44.56 d, the latter in the habitable zone (Crossfield et al. 2015). Even with more than 300 HARPS-N (High Accuracy Radial velocity Planet Searcher in North hemisphere) and HARPS high-precision radial velocities, it was not possible to determine the masses/densities of the two outer planets with a precision better than $3\sigma$ (Damasso et al. 2018; see also Kosiarek et al. 2019).
@@
Kepler-10 is a “historical” planetary system containing two transiting planets: Kepler-10 b, the first rocky planet discovered by the Kepler space mission (Batalha et al. 2011), with a radius of $R_{\rm b}=1.47\pm 0.03$ $\mathrm{R}_{\oplus}$ and an ultra-short orbital period (USP) of $P_{\rm b}=0.837$ d; and Kepler-10 c, which has both a larger radius $R_{\rm b}=2.35_{-0.04}^{+0.09}$ $\mathrm{R}_{\oplus}$ and a much longer orbital period $P_{\rm c}=45.294$ d (Fressin et al. 2011; Dumusque et al. 2014, hereafter D14). The planets orbit a 10.5-Gyr old GV star, which is thus expected to be magnetically quiet (average CaII activity index $\log{R^{{}^{\prime}}_{\rm HK}}\simeq-4.97$), and is located at a distance of $186$ pc.
@@
While there has been consensus on the mass and density of the inner rocky planet Kepler-10 b, determinations of the mass and density of Kepler-10 c in the literature have been largely discrepant, illustrating that the characterization of small planets with long orbital periods may be considerably problematic, even at low levels of stellar magnetic activity. For instance, with 148 high-precision RVs collected with the HARPS-N spectrograph at the Telescopio Nazionale Galileo (Cosentino et al. 2012, 2014), D14 found a RV semiamplitude of $K_{\rm c}=3.26\pm 0.36$  m s-1, which implies a planetary mass and density of $M_{\rm c}=17.2\pm 1.9$ $\mathrm{M}_{\oplus}$ and $\rho_{\rm c}=7.1\pm 1.0\leavevmode\nobreak\ \rm g\,cm^{-3}$. By using more sparse 72 RV measurements acquired with the HIRES (HIgh Resolution Echelle Spectrometer) spectrograph at the Keck telescope (Howard et al. 2010), Weiss et al. (2016, hereafter W16) found $K_{\rm c}=1.09\pm 0.58$  m s-1, which corresponds to $M_{\rm c}=5.7_{-2.9}^{+3.2}$ $\mathrm{M}_{\oplus}$ and $\rho_{\rm c}=2.4_{-1.2}^{+1.4}\leavevmode\nobreak\ \rm g\,cm^{-3}$ (one may actually even wonder whether, with a statistical significance of 1.9 $\sigma$, the signal of Kepler-10 c is truly detected in the HIRES RV data). By combining the 148 HARPS-N and the 72 HIRES RVs, W16 obtained $K_{\rm c}=2.67\pm 0.34$  m s-1, and hence $M_{\rm c}=13.98\pm 1.79$ $\mathrm{M}_{\oplus}$ and $\rho_{\rm c}=5.94\pm 0.75\leavevmode\nobreak\ \rm g\,cm^{-3}$. However, this combination of HARPS-N and HIRES data did not solve the issue of the evident mass discrepancy (see Fig. 5, bottom right panel in W16), but led to an average mass, which was found to be closer to the value found by D14 because the number of HARPS-N RVs was approximately twice that gathered with HIRES.
@@
Table 1: Radial-velocity semiamplitudes ($K$) and corresponding masses ($M_{\rm p}$) of the Kepler-10 planets from the literature.
@@
In addition to comparing the mass determinations obtained with HARPS-N and HIRES, W16 confirmed the variations in the Kepler-10 c transit times with an amplitude of $\sim 5$ min, which were previously discovered by Kipping et al. (2015). Such transit timing variations (TTVs) indicate the presence of an additional planet in the system, Kepler-10 d, which dynamically perturbs Kepler-10 c. In an attempt to take the constraints of TTVs into account, W16 performed a preliminary combined fit of TTVs and RVs, though (i) considering for simplicity a limited number of possible orbital configurations, near 2:1 or 3:2 resonance, for Kepler-10 d, and (ii) fixing the so-called TTV “super-period” at the peak of the periodogram of the TTVs at 475.22 d (see Sect. 6 in W16). By doing so, W16 found as their best solution that Kepler-10 d is in apparent 2:1 resonance with Kepler-10 c, and has a period of $101.36$ d and a mass of 6.84 $\mathrm{M}_{\oplus}$
@@
To try and reconcile the discrepancy in the RV semiamplitude, and hence planetary mass/density of Kepler-10 c, with the HIRES and HARPS-N data, Rajpaul et al. (2017) employed a framework based on Gaussian process (GP) regression with a quasi-periodic kernel to account for possible correlated noise (e.g., Haywood et al. 2014; Grunblatt et al. 2015), by simultaneously modeling the RVs and some activity indicators, such as the $\log{R^{{}^{\prime}}_{\rm HK}}$ and bisector activity indices (Rajpaul et al. 2015). As a result, they found $K_{\rm c}=1.27^{+0.42}_{-0.35}$  m s-1and $K_{\rm c}=1.64^{+0.42}_{-0.34}$  m s-1from the HIRES and HARPS-N data, respectively, and $K_{\rm c}=1.41^{+0.25}_{-0.23}$  m s-1when combining both datasets. The latter $K_{\rm c}$ would imply $M_{\rm c}=7.37^{+1.32}_{-1.19}$ $\mathrm{M}_{\oplus}$ and $\rho_{\rm c}=3.14^{+0.63}_{-0.55}\leavevmode\nobreak\ \rm g\,cm^{-3}$, which are considerably lower than those measured by both D14 and W16. From model comparison using Bayesian evidence, Rajpaul et al. (2017) also found that the three-planet model was favored over the two-planet one, and reported both the orbital period and mass of Kepler-10 d, namely $P_{\rm d}=102\pm 1$ d and $M_{\rm d}=5.90^{+1.70}_{-1.01}$ $\mathrm{M}_{\oplus}$ , in agreement with the findings of W16. Moreover, they found a higher evidence for the GP model, and derived a period of $55.5\pm 0.8$ d for the quasi-periodic variations, though it is not clear whether that period is related to the stellar rotation (Rajpaul et al. 2017).
@@
Table 1 summarizes the RV semiamplitudes and corresponding masses of the Kepler-10 planets as retrieved in the different aforementioned works from the literature. Since it is clear from all these works that Kepler-10 is a complex planetary system, we continued to monitor it in 2017, 2019, and 2020, by practically doubling the number of HARPS-N RVs within the HARPS-N/GTO program. After extracting most of the HARPS-N RVs with upgraded spectra reduction softwares (Sect. 2.2), we analyzed them with several tools (Sects. 3.3 and 3.4), also in combination with the Kepler-10 c TTVs (Sect. 3.6). In this way we improved the characterization of the Kepler-10 planetary system (Sect. 4) with regard to both its orbital architecture and the determination of precise masses, and thus possible compositions, of the Kepler-10 b and Kepler-10 c transiting planets. This in turn allowed us to put additional constraints on the formation and evolution of the Kepler-10 planetary system (Sect. 5).
@@ 2.1 Kepler photometry
Even though the Kepler light curve was analyzed in several previous works (Batalha et al. 2011; Fogtmann-Schulz et al. 2014; D14; W16; Dai et al. 2019; Singh et al. 2022), we performed a new analysis of the Kepler data to (i) re-compute the twenty-four transit times of Kepler-10 c in short (58 s) cadence, and the first two transit epochs observed in long-cadence mode (29.4 min), which are not provided by W16; and (ii) re-derive the Kepler-10 c transit parameters.
@@ 2.2.1 HARPS-N radial velocities
In total, we gathered 308 HARPS-N spectra, 55 of which before the failure of the red side of the charge-coupled device (CCD) in late September 2012 (Bonomo et al. 2014), and 253 after the replacement of the CCD. The former 55 spectra were reduced with the original HARPS-N data reduction software (DRS) version 3.7, and the latter 253 with the ESPRESSO (Echelle SPectrograph for Rocky Exoplanets and Stable Spectroscopic Observations) DRS version 2.3.5 adapted to HARPS-N (Dumusque 2021). The radial velocities were extracted from all the spectra by cross-correlating them with a G2V stellar template (e.g., Pepe et al. 2002). The new DRS pipeline allows for better long-term RV accuracy, due to a careful selection of non-saturated thorium and argon lines that are used to calibrate the spectrograph wavelength solution. In the DRS-3.7 Kepler-10 RV data, significant variations in the flux of the original thorium-argon calibration lamp induced an artificial quadratic long-term trend of $\sim 5$  m s-1 until the replacement of the thorium-argon lamp at the beginning of June 2020, which further produced an RV offset at the level of a  m s-1. Those systematics disappear or are at least strongly mitigated to a level below the  m s-1  when using the new pipeline. In addition, the new DRS uses a novel algorithm to determine the wavelength solution that is more stable from calibration to calibration. The night-to-night RV rms offset due to wavelength solution calibration is reduced from $\sim$80 to $\sim$50 cm s-1(Dumusque 2021). However, no significant improvements from the new DRS are expected for the first 55 HARPS-N spectra taken with the old CCD, given both the relatively small amount of data and the short timespan coverage. For this reason, we used the original HARPS-N DRS (version 3.7) for those spectra.
@@
Given the high complexity of the Kepler-10 system, as shown by the discrepancy in the Kepler-10 c mass determination from previous works, we also made use of the YARARA-v2 tool (Cretignier et al. 2021, 2022) to correct for telluric lines and subtle instrumental effects that cannot be accounted for by the new DRS pipeline, as well as for possible stellar activity variations, despite the low stellar activity level. In doing so, we had to discard 17 out of 253 HARPS-N spectra ($\sim 7$%) because they did not pass the signal-to-noise ratio (S/N) and vetting criteria of YARARA-v2.
@@
Both the 55 DRS-3.7 and 236 YARARA-v2 HARPS-N RVs used in this work are shown in Fig. 1. They were released by Bonomo et al. (2023) along with the stellar activity indicators full width at half maximum (FWHM), contrast and bisector span of the cross-correlation function (CCF), and the CaII H&K S-index and $\log{R^{{}^{\prime}}_{\rm HK}}$ [^1] We point out that the CCF contrast and FWHM activity indicators are affected by strong variations, which are highly correlated, due to changes of the HARPS-N focus. However, this did not result in any systematic RV variation, as the product $\rm Contrast\cdot FWHM$ is conserved.. The improvement achieved with YARARA-v2 can be seen in Fig. 2, which shows the increase in power of the peaks at the expected orbital periods of both Kepler-10 b and Kepler-10 c (from Kepler photometry) in the generalized Lomb-Scargle (GLS, Zechmeister & Kürster 2009) periodogram of the 236 HARPS-N RVs reduced with the new DRS+YARARA-v2 (bottom panel) compared to the new DRS (top panel). For this reason, we used the former for the 236 RVs collected from 2013 to 2021 (no remarkable improvement is instead expected for the 55 RVs gathered earlier for the same aforementioned reasons that we did not apply the new DRS for their extraction). We note that we checked for possible outliers in the HARPS-N RVs by using Chauvenet’s criterion (e.g., Bonomo et al. 2023), and found none.
@@
Figure 1: Kepler-10 radial velocities. Filled blue and magenta circles show the HARPS-N data collected with the new and old CCD, respectively, and empty green triangles display the HIRES measurements. Radial-velocity zero points as determined with the DE-MCMC analysis (Sect. 3.4.1 and Table LABEL:table_harpsn-hires_param) were subtracted from each dataset.
@@
Figure 2: Generalized Lomb-Scargle periodograms of the 236 HARPS-N radial velocities as reduced with the new DRS (top panel) and the new DRS+YARARA-v2 (bottom panel). The power of the periodograms was normalized by the 1% false alarm probability (FAP) level. Note the increase in power of the peaks at the periods of Kepler-10 b and Kepler-10 c (vertical green lines) with the YARARA-v2 reduction.